NvidiaIntely (nvidiaintely):

On a vacation, DeVante and his family traveled 382 miles at 64 mph. How long did the trip take? (Round to the nearest hundredth) 0.168 hours 24,448 hours 5.69 hours 5.97 hours

NvidiaIntely (nvidiaintely):

@Kevin

OpenStudy (kevin):

v = s/t 64 = 382/x x = 5.97

NvidiaIntely (nvidiaintely):

thanks

OpenStudy (kevin):

yw